I need help solving this quadratic equation by factoring:
t(3t+1)=2

Found 2 solutions by Alan3354, nerdybill:
Answer by Alan3354(69443)

I need help solving this quadratic equation by factoring:
t(3t+1)=2
----------
t(3t+1)=2
3t^2 + t = 2
3t^2 + t - 2 = 0 Always get it into this form. Well, almost always.
(t + 1)*(3t - 2) = 0
Either (t+1) = 0, or
(3t-2) = 0
-------------
t = -1
or t = 2/3

Answer by nerdybill(7384)

t(3t+1)=2
expanding the left:
3t^2 + t = 2
moving everything to the left, by subtracting 2 from both sides:
3t^2 + t - 2 = 0
Now, factor:
(3t-2)(t+1) = 0
.
Set each term to zero:
3t-2 = 0
3t = 2
t = 2/3
.
t+1 = 0
t = -1
.
Therefore, your solution set is:
t = {-1, 2/3}